Find the limits of the following sequences, if they exist. n+ (-1) COS n An = an = n
Answer #1

Solution 0 an= h+ch h . an = 1 + cu an = loimo o [1+(azh] win = I+D ☺ an = cosin Loing mana lorem non cosa = Ox oscillating e